OpenStudy (jesuslover11):

Which of the following is a true statement? a) it is possible for two rectangles to have the same area, but only if they also have the same perimeter b) it is possible for two rectangles to have the same area without having the same perimeter c) it is possible for two squares to have the same area without having the same perimeter d) it is possible for two squares to have the same perimeter without having the same area

OpenStudy (rational):

consider two rectangles : 3x2 6x1

OpenStudy (rational):

do they have same area ?

OpenStudy (jesuslover11):

yes

OpenStudy (rational):

good, what about perimeter ?

OpenStudy (jesuslover11):

can you give me an example of perimeter

OpenStudy (rational):

3x2 Area = 3*2 = 6 Perimeter = 2(3+2) = `10` 6x1 Area = 6*1 = 6 Perimeter = 2(6+1) = `14`

OpenStudy (rational):

Clearly both rectangles have same area but different perimeters

OpenStudy (jesuslover11):

so b

OpenStudy (rational):

Yep!

OpenStudy (jesuslover11):

Thank you so much!!!

OpenStudy (rational):

yw